Show that C (n+1, k) = C (n, k -1) + C (n, k)


1
Expert's answer
2021-10-25T15:29:26-0400
"C(n+1,k)=\dbinom{n+1}{k}=\dfrac{(n+1)!}{k!(n+1-k)!}"

"C(n,k-1)=\dbinom{n}{k-1}=\dfrac{n!}{(k-1)!(n-k+1)!}"

"C(n,k)=\dbinom{n}{k}=\dfrac{n!}{k!(n-k)!}"

Then


"C(n,k-1)+C(n,k)"

"=\dfrac{n!}{(k-1)!(n-k+1)!}+\dfrac{n!}{k!(n-k)!}"

"=\dfrac{n!(k+n-k+1)}{k!(n-k+1)!}=\dfrac{n!(n+1)}{k!(n-k+1)!}"

"=\dfrac{(n+1)!}{k!(n+1-k)!}=C(n+1,k)"